Tr3:

Which car traveled at the slowest rate? A) 300 miles in 4 hours Eliminate B) 450 miles in 6 hours C) 750 miles in 10 hours D) 800 miles in 12 hours

Tr3:

@Felicity96 @AnimeGhoul8863

Shadow:

\[\frac{ 300mi }{ 4h} , \frac{ 450mi }{ 6h }, \frac{ 750mi }{ 10h }, \frac{ 800mi }{ 12h }\] Divide, in order to get mph (miles per hour). The lowest number is the lowest rate of miles per hour.

Shadow:

@Tr3 Do you need help?

Tr3:

NAH caps